please explain the concept of slope

Grade:10

1 Answers

Simran Bhatia
13 years ago

Dear Teja,

Consider a function as :  y = f(x).

now if you find the rate of range of this particular function with respect to x:

i.e   dy/dx ; then this rate of change is known as slope.

In more simpler words consider follwing values:

at x = 3     :   y = 7

at x = 5     :   y = 10

So change in y   = 10 - 7 = 3.

and change in x  = 5 -3 = 2.

So slope can also be defined as :    change in y / change in x

                                                   =   3/2 = 1.5.

I hope now you have understood the concept of slope.

askIITians.com provides online iit jee courses and IIT JEE Test Series with IITians. Click here to get free online test series and check your status timely or you can join us as our registered user for getting best iit jee study material or iit jee test series.

Click here to find IIT JEE Packages or course details with askIITians.

for further details visit the folowing links:

http://www.askiitians.com//iit-jee-online-coaching.aspx

http://www.askiitians.com/iit-jee-test-series.aspx

http://www.askiitians.com/iit-study-material/

Askiitians provide you all the best possible information, feel free to ask.

all the best.

thanks and regards.

Akhilesh Shukla